The solubility of solid nickel hydroxide, Ni(OH)2, is governed by its Ksp= 6x10^-16. a) Write the equation for this dissolution reaction, and the equilibrium expression. b) Nickel ions undergo three complexations with hydroxide in basic solutions. Assume OH concentration to be 0.0010 M and pH to be 11, calculate the concentrations of Ni2+, NiOH+, Ni(OH)2, and Ni(OH)3-.

logK1= 4.1

logK2= 3.9

logK3= 3
Answer #1

② The solubility product of Nicola is Kap= 6*10-16 ay The dissolution equation for solid Në (OH), is given as Ni (OH) (3) - Nconcentration of Ni COH)* :- The completation reaction is - Nihag & OH Cag) Ne (OM) * eas) The equilibrium expression is m. R- [NE (0*.] a kex [Nier] [ou] Here.ly, a 3-9; R = 103.9 2 7-9x108 . [Ni (OH)] = 7.9 x 1081-5x10-10 (2+0.001) m 24.74 X10-12 M